OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)? -x + y = 2 -x – y = 2 x – y = 2 -x – y = 0 x + y = 0

Vocaloid (vocaloid):

1. find the slope of the original equation 2. take the negative reciprocal of the slope obtained in step 1 3. write an equation in the form y = mx + b, using the slope obtained in step 2 and the coordinates (-1,-1) 4. rewrite the equation in standard form

OpenStudy (anonymous):

can u walk me through each step

OpenStudy (anonymous):

Here you go

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

what steps are you confused about?

Vocaloid (vocaloid):

1. find the slope of the original equation -x + y = 7 y = x + 7, so the original slope is 1 2. take the negative reciprocal of the slope obtained in step 1 the negative reciprocal of the original slope is -1 3. write an equation in the form y = mx + b, using the slope obtained in step 2 and the coordinates (-1,-1) y = mx + b -1 = (-1)(-1) + b b = -2 so y = -x -2 4. rewrite the equation in standard form y = -x -2 -x - y = 2

OpenStudy (anonymous):

A plot of the lines is attached